OpenStudy (anonymous):

How to factor 2x^2 – 6x – 8.? I was trying to solve it the modern way but I was told to solve it by finding GCF and then making 2x^2 a x^2. The problem is, why does the equation "multiply first and last coefficient" and find sum for middle coefficient not work for this?

OpenStudy (zzr0ck3r):

factor a 2 out 2(x^2-3x-4) 1*-4 = -4 1+-4 = -3 2(x-4)(x+1)

OpenStudy (zzr0ck3r):

you could have use 2*-8 = 16 2+-8 = -6 so 2x^2-6x-8 = 2x^2+2x-8x-8 = 2x(x+1)+-8(x+1) = (2x-8)(x+1) = 2(x-4)(x+1)

OpenStudy (zzr0ck3r):

that should say 2*-8 = -16